# Lightning declines over shipping lanes following regulation of fuel sulfur emissions

### v2

@@ Acknowledgements
SST-driven fluxes on heat and moisture may also be important for setting the stage for convection to occur and for driving instability; however, the connection between SSTs and lightning is less obvious than that for the more direct measures of instability (CAPE and precipitation) used here. For a more in-depth analysis of the SST-pattern drivers of convection and ITCZ migration in this region, we refer the reader to Zhang et al zhang_how_2022.
@@
We use the GPM precipitation feature database from Chuntao Liu’s group to probe the shipping lane enhancement of cold, strong reflectivity echoes liu_cloud_2008. Note that this approach to examining reflectively is slightly different from that done in Thornton et al. Thornton et al., (2017), but follows that in Blossey et al Blossey et al., (2018). Due to infrequent sampling and small swath width, the reflectivity signal shown here requires very long records to show the effect of the shipping lane (previous studies, such as Blossey et al Blossey et al., (2018), use the full TRMM record, which is well over a decade). Splitting the relatively short (8 year) GPM record of reflectivity into a pre- and post-IMO period therefore greatly reduces the statistical power of the retrievals. Indeed, the peak in reflectivity established over the shipping lane between 2016 and 2019 has diminished (Figure S5). However, the difference between the two periods is not statistically significant (p less than 0.05), and we therefore refrain from concluding that there has or has not been a change to reflectivity above the shipping lanes.
@@
While a correspondence between aerosol optical depth (AOD) and CCN has been documented in some highly polluted regions, this relationship is not established for the typically clean (AOD<0.02) regions of interest here. Therefore, as noted in the main manuscript, we expect that perturbations to CCN in the shipping lane regions is best represented by MODIS Nd retrievals, rather than by AOD because 1) clear sky conditions are rare in this region leading to limited AOD retrievals and consequently limited statistical power, 2) the background AOD is very low, such that an expected perturbation is barely above the limit of detection, and 3) AOD being proportional to aerosol cross sectional area columns does not directly correspond to CCN number at cloud base, the subject of this paper. For example, as shown in Figure S5, AOD estimates from MERRA-2 reanalysis, nudged by both their aerosol model and the limited AOD observations by MODIS, exhibit no observable enhancement over the shipping lanes prior to or after the 2020 IMO regulation. This lack of AOD enhancement is somewhat expected given that ship emissions of aerosol particles are concentrated in smaller, less radiatively active particle sizes [see, e.g., Hobbs et al., Hobbs et al., (2000) and Seppala et al. Seppälä et al., (2021)], statistical sampling is limited by infrequent cloud-free scenes, and that particle scavenging and dilution rates in these regions are relatively high due frequent convection.

### v3

@@ Acknowledgements
SST-driven fluxes on heat and moisture may also be important for setting the stage for convection to occur and for driving instability; however, the connection between SSTs and lightning is less obvious than that for the more direct measures of instability (CAPE and precipitation) used here. For a more in-depth analysis of the SST-pattern drivers of convection and ITCZ migration in this region, we refer the reader to Zhang et al (2022) (doi.org/10.1175/JCLI-D-21-0870.1).
@@
We use the GPM precipitation feature database from Chuntao Liu’s group to probe the shipping lane enhancement of cold, strong reflectivity echoes (doi.org/10.1175/2008JAMC1890.1). Note that this approach to examining reflectively is slightly different from that done in Thornton et al., (2017), but follows that in Blossey et al., (2018). Due to infrequent sampling and small swath width, the reflectivity signal shown here requires very long records to show the effect of the shipping lane (previous studies, such as Blossey et al., (2018), use the full TRMM record, which is well over a decade). Splitting the relatively short (8 year) GPM record of reflectivity into a pre- and post-IMO period therefore greatly reduces the statistical power of the retrievals. Indeed, the peak in reflectivity established over the shipping lane between 2016 and 2019 has diminished (Figure S5). However, the difference between the two periods is not statistically significant (p less than 0.05), and we therefore refrain from concluding that there has or has not been a change to reflectivity above the shipping lanes.
@@
While a correspondence between aerosol optical depth (AOD) and CCN has been documented in some highly polluted regions, this relationship is not established for the typically clean (AOD<0.02) regions of interest here. Therefore, as noted in the main manuscript, we expect that perturbations to CCN in the shipping lane regions is best represented by MODIS Nd retrievals, rather than by AOD because 1) clear sky conditions are rare in this region leading to limited AOD retrievals and consequently limited statistical power, 2) the background AOD is very low, such that an expected perturbation is barely above the limit of detection, and 3) AOD being proportional to aerosol cross sectional area columns does not directly correspond to CCN number at cloud base, the subject of this paper. For example, as shown in Figure S5, AOD estimates from MERRA-2 reanalysis, nudged by both their aerosol model and the limited AOD observations by MODIS, exhibit no observable enhancement over the shipping lanes prior to or after the 2020 IMO regulation. This lack of AOD enhancement is somewhat expected given that ship emissions of aerosol particles are concentrated in smaller, less radiatively active particle sizes [see, e.g., Hobbs et al., (2000) and Seppälä et al., (2021)], statistical sampling is limited by infrequent cloud-free scenes, and that particle scavenging and dilution rates in these regions are relatively high due frequent convection.